# SonarSplat: Novel View Synthesis of Imaging Sonar via Gaussian Splatting

### v1

@@ Abstract
In this paper, we present SonarSplat, a novel Gaussian splatting framework for imaging sonar that demonstrates realistic novel view synthesis and models acoustic streaking phenomena. Our method represents the scene as a set of 3D Gaussians with acoustic reflectance and saturation properties. We develop a novel method to efficiently rasterize learned Gaussians to produce a range/azimuth image that is faithful to the acoustic image formation model of imaging sonar. In particular, we develop a novel approach to model azimuth streaking in a Gaussian splatting framework. We evaluate SonarSplat using real-world datasets of sonar images collected from an underwater robotic platform in a controlled test tank and in a real-world river environment. Compared to the state-of-the-art, SonarSplat offers improved image synthesis capabilities (+2.5 dB PSNR). We also demonstrate that SonarSplat can be leveraged for azimuth streak removal and 3D scene reconstruction.
1 Introduction
Figure 1: We present SonarSplat, a method that uses posed sonar images to produce a novel 3D Gaussian Splat representation that enables novel view synthesis of realistic sonar images and coarse 3D reconstruction. With SonarSplat, the scene is represented with a set of 3D Gaussians that model the reflectivity and azimuth streaking probabilities of objects in the rendering process. This novel formulation enables additional features of SonarSplat, such as azimuth streak removal, enabling the recovery of clean sonar images from saturated ones.
@@ 1 Introduction
Acoustic sensors, such as imaging sonar, are commonly used for infrastructure inspection, large-area mapping, and target detection in underwater environments [12, 34, 20]. Unlike optical sensors, which are severely range-limited due to water column effects on light propagation, acoustic sensors can capture data at long ranges to provide critical information about subsea environments. Although sonar exhibits many desirable qualities such as longer range, invariance to lighting conditions, and the ability to discern material properties, there exist acoustic phenomena, including elevation ambiguity, azimuth streaking, and multi-path reflections, which make sonar interpretation difficult for operators and computer vision algorithms alike. Furthermore, the severe lack of large, public datasets for sonar slows progress in traditional computer vision tasks like object detection, segmentation, and reconstruction.
@@
Recently, neural radiance fields (NeRFs) [22] have demonstrated the potential for high-fidelity data synthesis and denoising of optical imagery [27, 6]. The benefits of neural rendering have been recognized by the underwater perception community, with prior work exploring neural rendering for 3D object reconstruction using underwater cameras [15, 35] and sonar data [30, 31, 44, 7, 43]. While these innovations are promising, training and processing NeRFs is extremely costly and time-consuming, making deploying them in real-time or on resource-constrained devices difficult. More recently, 3D Gaussian splatting (3DGS) [13] was developed as a faster alternative to NeRF. 3DGS utilizes sparse points in the form of 3D Gaussians to represent the scene. This allows 3DGS to preserve properties of continuous volumetric radiance fields while lowering computational costs. Gaussian splatting has been leveraged for underwater imagery [45, 16, 23]. Most relevant to our proposed work, ZSplat proposes a Gaussian splatting framework for RGB-sonar fusion [32]. However, ZSplat leverages the fusion of sonar data to improve the rendering of RGB images and does not focus on enabling high-fidelity data synthesis for sonar imagery or provide evaluation for the quality of rendered sonar images. Thus, there is a clear gap in the literature for a framework capable of efficient and effective sonar image synthesis.
In this paper, we propose SonarSplat (Fig. 1), a novel Gaussian splatting framework for imaging sonar that enables efficient and high quality novel view synthesis for imaging sonar data. We focus on modeling characteristic acoustic noise properties, specifically, azimuth streaking, or receiver saturation, which can make image interpretation and 3D reconstruction difficult. By modeling realistic noise effects, we demonstrate that we can leverage SonarSplat to perform efficient, high quality sonar image synthesis and noise removal. Lastly, we also demonstrate SonarSplat’s ability to learn geometry of the scene, overcoming the elevation ambiguity inherent in imaging sonar image formation. Our main contributions are summarized as follows:
@@
Code and data will be made publicly available on a project website after a double-blind review.
2 Related Work
@@
Figure 2: Overview of SonarSplat. Our method takes as input a sensor pose and an initial set of 3D Gaussians representing the scene. Then, we transform the 3D Gaussians into the sensor’s image space. We splat the reflectivity parameter $\nu_{k}$ to get the unsaturated image $I_{u}$. Additionally, we optimize and splat per-Gaussian azimuth streaking probabilities $p_{k}$. All the probabilities in range interval $r_{i}$ are considered in our novel Adaptive Gain Module, which adjusts the receiver gain applied to $I_{u}$. Finally, we produce $\hat{I}$ by multiplying the gain $A$ by $I_{u}$. All parameters are optimized using gradient descent by taking losses with respect to the sonar image pixel values. All sonar images shown are polar (range-azimuth) coordinates.
2.1 Radiance Field Methods
@@
NeRFs represent scenes implicitly using neural networks and have been explored for a variety of applications, including novel view synthesis, inverse rendering, data generation, and image denoising [21, 8, 27, 6]. Later works improved NeRFs through improved sampling of the volumetric representation and improvements in rendering speed with alternative encoding methods [24, 2]. Despite their vast success as a learned scene representation, a key struggle for the practical application of NeRFs is the long training time and the slow rendering speeds, primarily due to the computationally expensive volumetric rendering process [21]. A recent development in radiance field methods is Gaussian splatting, which uses rasterization instead of volumetric rendering to optimize and produce a scene representation [13]. The resulting renders from 3DGS are both high in quality and reach unprecedented rendering speeds, enabling fast rendering for various real-time applications [19, 18].
@@
Despite their impressive performance and rapid adoption, radiance field methods were developed with optical sensors in mind, limiting their direct application to non-visual sensors. Recent developments have integrated the physics-based image formation models of various non-visual sensors, including radar, LIDAR, sonar, thermal cameras, and event cameras into NeRF and Gaussian splatting frameworks for various vision tasks [3, 4, 10, 30, 17]. In this work, we propose a novel Gaussian splatting framework for imaging sonar for underwater applications.
2.2 Perception Challenges for Imaging Sonar
@@
Imaging sonar is a popular underwater sensor for perception tasks like object detection, segmentation, reconstruction, and inspection [30, 9, 12, 38, 25]. The key challenges with working with imaging sonars for these tasks are properties inherent with the acoustic image formation model [29]. The sensor is susceptible to multi-path effects, highly nonlinear noise patterns, and phenomena like azimuth streaking [28]. Azimuth streaking, in particular, can make sonar image understanding difficult by suppressing returns from certain azimuth angles, causing black bars in the sonar image [28]. In addition, the sensor inherently loses information from 3D into 2D when capturing the acoustic returns, causing an effect known as elevation ambiguity [40]. This ambiguity has made the successful application of imaging sonar to 3D vision tasks difficult [25, 11, 42]. Yet, imaging sonars still remain an extremely valuable sensor for underwater imaging, especially in long range imaging and deep water applications where optical sensors may fail [39].
Given its importance to underwater perception, prior work has proposed sonar simulators to validate perception algorithms without the cost, time commitment, and difficulty of real-world data collection [29, 5]. Although these works model complex interactions with the environment, such as multi-path reflections and azimuth streaking, there still exists a sim-to-real gap inherent with simulators. Furthermore, simulators also require the manual tuning of acoustic and scene parameters to match the distribution of data desired by the end-user. Rather, our proposed work proposes a Gaussian splatting framework that can learn from real data to produce realistic sonar data from novel poses, which may be useful for augmenting datasets and addressing the data scarcity problem.
2.3 Novel View Synthesis for Imaging Sonar
@@
Prior work has incorporated the sonar image formation model into signed distance field (SDF) and NeRF formulations [30, 7]. Neusis [30] presents a NeRF framework for dense 3D reconstruction of objects using an imaging sonar, specifically, a forward-looking sonar (FLS). Neusis proposes a differentiable volumetric renderer to model the propagation of acoustic waves to synthesize imaging sonar measurements. By using this implicit representation and multiple sensor measurements, Neusis attempts to resolve the elevation ambiguity present in the image data to construct a 3D mesh of the environment. Neusis-NGP [44] is an extension of Neusis that leverages multi-resolution hash encodings to produce a bathymetric heightmap of the environment. To perform novel view synthesis with Neusis and Neusis-NGP, every 2D pixel that makes up a sonar image is sampled along its elevation ambiguity and then evaluated by the rendering networks to determine the intensity of each point. These points are then summed to extract the return at the specific pixel coordinate. We show that this process is time-consuming and does not produce photorealistic sonar imagery. Differentiable Space Carving (DSC) [7] proposes a faster method that models echo probabilities to carve out a mesh of the scene. This method is similar to Neusis, but avoids having to implicitly represent the scene as intensities. While this method is faster, it still requires sampling points at every pixel and then evaluating the occupancy to perform sonar novel view synthesis. Through our experiments, we show that DSC is also unable to produce high-quality novel view synthesis of imaging sonar.
@@
Following recent developments in 3DGS, Z-Splat [32] introduces a novel method for fusing camera and sonar data, enhancing RGB image rendering by incorporating depth supervision from sonar returns. To the best of our knowledge, Z-Splat [32] is the only Gaussian splatting paper that incorporates acoustic sensor data, with demonstrations on underwater data. This is done by splatting Gaussians onto the X/Z plane and calculating a depth loss between the sonar returns and the splatted Gaussians. While this process improves the placement of Gaussians for better camera renderings, Z-Splat oversimplifies the splatting process by not accounting for any sonar-specific phenomena. More importantly, Z-Splat only splats the opacities of Gaussians, and does not consider material properties like acoustic reflectance. This leads to poor sonar renderings, which limit its application to sonar novel view synthesis. We demonstrate that SonarSplat, using range-azimuth splatting, produces more accurate novel view synthesis and can also restore sonar images to remove azimuth streaking artifacts.
3 Technical Approach
@@
Fig. 2 shows an overview of SonarSplat, which uses sonar images and their corresponding sensor poses to optimize a scene representation. The scene is represented as a set of 3D Gaussian primitives, each parametrized with a mean, covariance, opacity, acoustic reflectivity and azimuth streaking probability. First, the Gaussians that fall within the view frustum of the sonar are rasterized into a range-azimuth image. Concurrently, we render an azimuth streak probability image using the per-Gaussian azimuth streak probabilities. Based on the relative probabilities in each range bin, we use a novel adaptive gain module to adjust the acoustic returns and capture azimuth streaking effects. Finally, we produce a sonar image with azimuth streaks that can be optimized by taking a loss with respect to the ground truth sonar image. We provide further detail of each of these components in the remainder of this section.
3.1 Gaussian Splatting Primer
This section provides background information about traditional Gaussian splatting for optical images. The standard 3D Gaussian splatting framework models the scene using a set of $N$ 3D Gaussian basis functions, $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$, each parametrized with a mean $\mu_{i}\in\mathbb{R}^{3}$, scaling $s_{i}\in\mathbb{R}^{3}$, quaternion $q_{i}\in\mathbb{R}^{4}$, opacity $o_{i}\in\mathbb{R}$, color $c_{i}\in\mathbb{R}^{3}$, encoded as spherical harmonics [13]. In this work, we augment this parametrization with acoustic-specific variables: acoustic reflectivity $\nu_{i}$ and azimuth streak probability $p_{i}$. The covariance matrix is constructed from the rotation, $q_{i}$, and scaling, $s_{i}$, of each Gaussian, each of which is converted to a rotation matrix, $R_{i}\in\mathbb{R}^{3\times 3}$, and scaling matrix, $S_{i}\in\mathbb{R}^{3\times 3}$ [13]:
$\Sigma_{i}=R_{i}S_{i}S_{i}^{T}R_{i}^{T}$ (1)
At a given camera pose $W\in SE(3)$, an image $I_{\mathcal{G}}$ can be rendered in a differentiable manner using a point-based rasterizer. The color, $C$, at a pixel, $p$, is computed by blending $\widehat{N}$ sorted Gaussians that overlap that pixel:
$C(p)=\sum_{i=1}^{\widehat{N}}c_{i}o_{i}T_{i}G_{i}$ (2)
where $T_{i}=\prod_{j=1}^{i-1}\left(1-o_{i}G_{i}\right)$ is the transmittance of the Gaussian $G_{i}$. The scene is optimized by computing the loss, $\mathcal{L}_{GS}$, between the rendered image $I_{\mathcal{G}}$ with the ground truth image $I_{\text{GT}}$, through summing $L_{1}$ loss and SSIM loss, $\mathcal{L}_{\text{SSIM}}$, as follows:
$\mathcal{L}_{GS}=\lVert I_{\mathcal{G}}(W)-I_{\text{GT}}(W)\rVert_{1}+\mathcal{L}_{\text{SSIM}}(I_{\mathcal{G}}(W),I_{\text{GT}}(W))$ (3)
3.2 Imaging Sonar Image Formation Model
@@
To adapt the optical Gaussian splatting framework to accommodate imaging sonar, we must incorporate the acoustic image formation model. Imaging sonars emit acoustic energy from a transmitter then listen for returns on the receiver. Imaging sonar is a time-of-flight sensor that can provide the echo intensities at a given range and azimuth bin. Importantly, imaging sonar is not able to distinguish the elevation angle of a return, leading to an elevation angle ambiguity. This ambiguity makes direct use of imaging sonar difficult for 3D reconstruction and triangulation tasks [41, 11].
@@
A sonar image is defined as $I\in\mathbb{R}^{N_{r}\times N_{a}}$ where $N_{r}$ denotes the number of range bins, and $N_{a}$ denotes the number of azimuth bins. In spherical coordinates, the sonar has a vertical field of view ($\phi_{min},\phi_{max}$), a horizontal field of view ($\theta_{min},\theta_{max}$), and a maximum range, $R_{max}$. Finally, taking $\frac{R_{max}}{N_{r}}$ gives us the range resolution, $\epsilon_{r}$, and taking $\frac{\theta_{max}-\theta{min}}{N_{a}}$ gives us the azimuthal resolution, $\epsilon_{a}$.
@@
where $E_{e}$ is the intensity of emitted sound, $r$ is the range of the point, $T(r,\theta,\phi)$ is the transmittance term, and $\sigma(r,\theta,\phi)$ is the density of the point. Note that this transmittance term is equivalent to the term used in volumetric rendering such that $T$ serves as the probability that a ray travels to point $(r,\theta,\phi)$ without hitting another particle.
3.3 Range/Azimuth Splatting Model
To evaluate this integral, prior methods use sampling and quadrature [30, 31, 7]. However, motivated by the increased efficiency of splatting techniques, we explore a point-based approach adapted to the sonar image formation model. To this end, we represent the scene with a set of 3D Gaussian primitives $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$. Each of these Gaussians are parametrized with a mean $\mu_{i}\in\mathbb{R}^{3}$, scale $s_{i}\in\mathbb{R}^{3}$, orientation $q_{i}\in\mathbb{R}^{4}$, opacity $o_{i}\in\mathbb{R}$, acoustic reflectivity $\nu_{i}\in\mathbb{R}$, and azimuth streak probability $p_{i}\in\mathbb{R}$:
$G_{i}=\{\mu_{i},s_{i},q_{i},o_{i},\nu_{i},p_{i}\}$ (5)
where the covariance is computed as shown in Eq. 1.
To render sonar images, we first transform the means and covariances of the Gaussians into spherical coordinates. For a given 3D Gaussian in the sensor local frame in Cartesian coordinates, $G_{k}=\{\mu_{k}=\langle x_{k},y_{k},z_{k}\rangle,\Sigma_{k}\}$, the transformation to spherical coordinates is as follows [14]:
$\mu_{s}=\begin{bmatrix}r_{k}\\ \theta_{k}\\ \phi_{k}\end{bmatrix}=\begin{bmatrix}||\mu_{k}||\\ \arctan(y_{k},x_{k})\\ \arctan(z_{k},\sqrt{x_{k}^{2}+z_{k}^{2}})\\ \end{bmatrix}$ (6)
The covariance $\Sigma_{k}$ is transformed by using a first-order linearization of the coordinate transformation, which the Jacobian describes:

### v2

@@ Abstract
In this paper, we present SonarSplat, a novel Gaussian splatting framework for imaging sonar that demonstrates realistic novel view synthesis and models acoustic streaking phenomena. Our method represents the scene as a set of 3D Gaussians with acoustic reflectance and saturation properties. We develop a novel method to efficiently rasterize Gaussians to produce a range/azimuth image that is faithful to the acoustic image formation model of imaging sonar. In particular, we develop a novel approach to model azimuth streaking in a Gaussian splatting framework. We evaluate SonarSplat using real-world datasets of sonar images collected from an underwater robotic platform in a controlled test tank and in a real-world river environment. Compared to the state-of-the-art, SonarSplat offers improved image synthesis capabilities (+3.2 dB PSNR) and more accurate 3D reconstruction (52% lower Chamfer Distance). We also demonstrate that SonarSplat can be leveraged for azimuth streak removal.
I Introduction
@@
Acoustic sensors, such as imaging sonar, are commonly used for infrastructure inspection, large-area mapping, and target detection in underwater environments [12, 39, 24]. Unlike optical sensors, which are severely range-limited due to water column effects on light propagation, acoustic sensors can capture data at long ranges to provide critical information about subsea environments. Although sonar exhibits many desirable qualities such as longer range, invariance to lighting conditions, and the ability to discern material properties, there exist acoustic phenomena, including elevation ambiguity, azimuth streaking, and multi-path reflections, which make sonar interpretation difficult for operators and computer vision algorithms alike.
Figure 1: We present SonarSplat, a novel 3D Gaussian splatting framework that enables novel view synthesis of realistic sonar images, image de-streaking, and 3D reconstruction. We model the acoustic reflectivity and azimuth streak probabilities for all Gaussians, yielding a 3D scene representation that can be queried for a variety of downstream tasks.
@@
Radiance field methods offer a promising approach to addressing these challenges. Neural radiance fields (NeRFs) have demonstrated the potential for high-fidelity data synthesis, denoising, and 3D reconstruction using optical imagery [26, 45, 30, 5]. The benefits of neural rendering have been recognized by the underwater perception community, with prior work exploring neural rendering for 3D object reconstruction using underwater cameras [18, 41] and sonar data [33, 34, 50, 6, 49]. While these innovations are promising, training and processing NeRFs is costly and time-consuming, making deploying them in real-time or on resource-constrained devices difficult. More recently, 3D Gaussian splatting (3DGS) [14] was developed as a faster alternative to NeRF. Gaussian splatting has been leveraged for underwater applications as well [51, 19, 27]. Most relevant to our proposed work, ZSplat proposes a Gaussian splatting framework for RGB-sonar fusion [35]. However, ZSplat leverages the fusion of sonar data to improve the rendering of RGB images and does not focus on enabling high-fidelity data synthesis for sonar imagery or provide evaluation for the quality of rendered sonar images or 3D reconstruction. Thus, there is a clear gap in the literature for a framework capable of efficient and effective sonar image synthesis and 3D reconstruction using Gaussian splatting.
In this paper, we propose SonarSplat (Fig. 1), a novel Gaussian splatting framework for imaging sonar that enables efficient and high quality novel view synthesis, sonar image de-streaking, and 3D reconstruction for underwater robotic applications. To the best of our knowledge, SonarSplat is the first sonar-only method that leverages Gaussian splatting to perform these tasks. Our main contributions are summarized as follows:
@@
  • We introduce a novel densification strategy called the Elevation Sampling Densification Strategy (ESDS) that places Gaussians on elevation arcs and yields better novel view synthesis and 3D reconstruction.
@@
Code and data will be made publicly available on a project website after review.
II Related Work
@@
Figure 2: Overview of SonarSplat. Our method takes as input a sensor pose and an initial set of 3D Gaussians representing the scene. Then, we transform the 3D Gaussians into the sensor’s image space. We splat the reflectivity parameter $\nu_{k}$ to get the unsaturated image $I_{u}$. Additionally, we optimize and splat per-Gaussian azimuth streaking probabilities $p_{k}$. All of the probabilities in a range bin $r_{i}$ are considered in our novel Adaptive Gain Module, which adjusts the receiver gain applied to $I_{u}$. Finally, we produce $\hat{I}$ by multiplying the gain $A$ by $I_{u}$. All parameters are optimized using gradient descent by taking losses with respect to the sonar image pixel values. All sonar images shown are polar (range/azimuth) coordinates.
II-A Radiance Field Methods
@@
NeRFs represent scenes implicitly using neural networks and have been explored for a variety of robotic applications, including novel view synthesis, data generation, Simultaneous Localization and Mapping (SLAM), manipulation, image dehazing, and trajectory optimization [26, 2, 11, 37, 43, 5, 7]. Despite their success as a scene representation, a key challenge for the practical application of NeRFs is the long training time and the slow rendering speeds, primarily due to the computationally expensive volumetric rendering process [26]. A recent development in radiance field methods is Gaussian splatting, which uses more efficient rasterization instead of volumetric rendering to optimize and produce a scene representation, enabling usage in a variety of real-time applications [14, 23, 22, 25].
@@
Despite their impressive performance and rapid adoption, radiance field methods were developed with optical sensors in mind, limiting their direct application to non-visual sensors. Recent developments have integrated the physics-based image formation models of various non-visual sensors, including radar, LIDAR, sonar, thermal cameras, and event cameras into NeRF and Gaussian splatting frameworks for various vision tasks [3, 4, 9, 33, 20]. We are inspired by the faster rendering speeds and high fidelity of Gaussian splatting methods and propose a novel Gaussian splatting framework for imaging sonar.
II-B Perception Challenges for Imaging Sonar
@@
Imaging sonar is a popular underwater sensor for perception tasks like object detection, segmentation, reconstruction, and inspection [33, 8, 12, 44, 28]. However, the sensor is susceptible to multi-path effects, nonlinear noise patterns, and phenomena like azimuth streaking [31]. Azimuth streaking, in particular, can make sonar image understanding difficult by suppressing returns from certain azimuth angles, causing black bars in the sonar image [31]. In addition, the sensor inherently loses information from 3D into 2D when capturing the acoustic returns, causing an effect known as elevation angle ambiguity [46]. This ambiguity has made the successful application of imaging sonar to 3D vision tasks challenging [28, 10, 48].
II-C Novel View Synthesis for Imaging Sonar
@@
Prior work has incorporated the sonar image formation model into signed distance field (SDF) and NeRF formulations [33, 6, 50]. Neusis presents a volumetric rendering framework for dense 3D reconstruction of objects using an imaging sonar [33]. Neusis-NGP is an extension of Neusis that leverages multi-resolution hash encodings to produce a bathymetric heightmap of the environment [50]. Differentiable Space Carving (DSC) proposes a faster method that models echo probabilities to carve out a mesh of the scene [6]. Through our experiments, we show that none of these method are able to perform realistic looking novel view synthesis for imaging sonar, limiting their application to data generation and augmentation. Furthermore, querying these methods is time consuming and limits their application to real-time methods.
@@
To the best of our knowledge, ZSplat is the only Gaussian splatting paper that incorporates imaging sonar data, with demonstrations on underwater data. However, ZSplat oversimplifies the splatting process by not accounting for any sonar-specific phenomena. More importantly, ZSplat only splats the opacities of Gaussians and does not consider material properties like acoustic reflectance. This leads to poor sonar renderings, which limits its application to sonar novel view synthesis. We demonstrate that SonarSplat, which more accurately models sonar image formation, produces more realistic novel view synthesis, restores sonar images to remove azimuth streaking artifacts, and produces 3D reconstructions under a unified framework.
III Technical Approach
@@
Figure 2 shows an overview of SonarSplat, which uses sonar images and their corresponding sensor poses to optimize a scene representation. The scene is represented as a set of 3D Gaussian primitives, each parametrized with a mean, covariance, opacity, acoustic reflectivity and azimuth streaking probability. First, the Gaussians that fall within the view frustum of the sonar are rasterized into a range/azimuth image. Concurrently, we render an azimuth streak probability image using the per-Gaussian azimuth streak probabilities. Based on the relative probabilities in each range bin, we use a novel adaptive gain module to adjust the acoustic returns and capture azimuth streaking effects. Finally, we produce a sonar image with azimuth streaks that can be optimized by taking a loss with respect to the ground truth sonar image. We provide further detail of each of these components in the remainder of this section.
III-A Imaging Sonar Formation Model
@@
To adapt the optical Gaussian splatting framework to accommodate imaging sonar, we must incorporate the acoustic image formation model. Imaging sonar is a time-of-flight sensor that can provide the echo intensities at a given range and azimuth bin. Importantly, imaging sonar is not able to distinguish the elevation angle of a return, leading to an elevation angle ambiguity. This ambiguity makes direct use of imaging sonar difficult for 3D reconstruction and triangulation tasks [47, 10].
@@
A sonar image is defined as $I\in\mathbb{R}^{N_{a}\times N_{r}}$ where $N_{a}$ denotes the number of azimuth bins and $N_{r}$ denotes the number of range bins. In spherical coordinates, the sonar has a vertical field of view ($\phi_{min},\phi_{max}$), a horizontal field of view ($\theta_{min},\theta_{max}$), and a maximum range, $R_{max}$. Finally, taking $\frac{R_{max}}{N_{r}}$ gives us the range resolution, $\epsilon_{r}$, and taking $\frac{\theta_{max}-\theta{min}}{N_{a}}$ gives us the azimuthal resolution, $\epsilon_{a}$.
@@
where $E_{e}$ is the intensity of emitted sound, $r$ is the range of the point, $T(r,\theta,\phi)$ is the transmittance term, and $\sigma(r,\theta,\phi)$ is the density of the point. Note that this transmittance term is equivalent to the term used in volumetric rendering, serving as the probability that a ray travels to point $(r,\theta,\phi)$ without hitting another particle.
III-B Range/Azimuth Splatting Model
To evaluate this integral, prior methods use sampling and quadrature [33, 34, 6]. However, motivated by the increased efficiency of splatting techniques, we explore a point-based approach adapted to the sonar image formation model. To this end, we represent the scene with a set of 3D Gaussian primitives $\mathcal{G}=\{G_{k}\}_{k=1}^{N}$. We augment this parametrization with acoustic-specific variables: acoustic reflectivity $\nu_{k}$ and azimuth streak probability $p_{k}$. Formally, these Gaussians are parametrized with a mean $\mu_{k}\in\mathbb{R}^{3}$, scale $s_{k}\in\mathbb{R}^{3}$, orientation $q_{k}\in\mathbb{H}$, opacity $o_{k}\in\mathbb{R}^{+}$, acoustic reflectivity $\nu_{k}\in\mathbb{R}^{+}$, and azimuth streak probability $p_{k}\in\mathbb{R}^{+}$:
$G_{k}=\{\mu_{k},s_{k},q_{k},o_{k},\nu_{k},p_{k}\}$ (2)
where the covariance $\Sigma_{k}$ is computed as in [14].
To render sonar images, we first transform the means and covariances of the Gaussians into spherical coordinates. Let $\mu_{k}=\langle x_{k},y_{k},z_{k}\rangle~{}\text{and}~{}\Sigma_{k}$ be the mean and covariance in the sensor’s frame, respectively. The transformation to spherical coordinates is as follows [15]:
$\mu_{s}=\begin{bmatrix}r_{k}\\ \theta_{k}\\ \phi_{k}\end{bmatrix}=\begin{bmatrix}||\mu_{k}||\\ \arctan(y_{k},x_{k})\\ \arctan(z_{k},\sqrt{x_{k}^{2}+y_{k}^{2}})\\ \end{bmatrix}$ (3)
The covariance matrix $\Sigma_{k}$ is transformed by using a first-order linearization of the coordinate transformation, which the Jacobian describes: